Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News, Travel, Visitors (McMullin)

  • Overall status: C config; intense observing pressure in the 1100-2000 LST range; correlator testing previews of coming functionality; initial staging for OPT next release. OST turnover on Wednesday - no impact in general but commissioning tests will need to be formally prioritized (assigned a rating). Reminder next week is a holiday week (Thu-Fri).
  • Key dates:
  • Staff Travel:
    • Vivek away until 02 Dec
  • Visitors:
    • Bob Sault

Commissioning priorities for the week (McMullin, Chandler)

Single maintenance day week:
  • Mon-Wed+ daytime (Recirculation testing/fast dumps/additional subreflector tests/data suck-out checks; OSRO 2011 tests)
  • night commissioning tests (flux density check; Ku SEFD)
  • Demo Science
  • Science (review Wednesday whether additional time Thu-Fri can be obtained)
  • Interrupt status for additional 3-bit testing; expected to be lab work only this week

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
  • WIDAR (Rupen):
    • Several changes last week (update to StB to enable new monitoring data (under review); CBE nodes updated to latest version of operating system (tested in a single node to vet). Issue with filter gains required a temporary revert in CMIB version to handle setting the default values properly; change that enables CM to set this will be coordinated later (issue that had caused problems with the narrow band observing briefly). Missing data noted (18 sec at the end of a scan prior to a band change that leads to a sideband switch; 10 sec at the beginning of a band change-being explored); mislabeled scans seems to be resolved; still seeing some missing scans - mostly isolated to last scan or scans immediately after OSRO. Static recirculation tests (a.k.a. hardware stacking - providing multiple baseline board pairs per subband to improve number of spectral channels) were successful (initial: 96 to cover 6 subbands=125 kHz channels over 128 MHz). Exploring sundry other issues: >4 GHz separation problem; EA09 delays; phase jumps on 12 B/D, 14 A/C, 27 A/C and B/D). Splatter tests (Sault).
  • Computing (Robnett):

Software status and updates (Butler)

Observing recommendations update (Ops)

Commissioning Reports

Demo science: Chandler/McMullin: Planning for C configuration and beyond

Calibration: Marvil: Gain curve polynomial fitting

Group reports

Polarization: Moellenbrock: Calibration strategy testing

Spectral Line: Momjian: Follow-up on blank field

In preparation

  • OSRO: Krauss: Uniform noise across sub-band boundaries.
  • * Flagging*: Venkata: Autoflagging in CASA (Coming soon)
  • Polarization: Myers: Multi-band leakage/PA
  • Array: Dhawan/Rupen: RF switch tests with rapid monitoring; issue still affecting; being replaced as identified. Looking at acquiring new switches to solve the problem in the long term.

Commissioning Test Planning

  • High Frequency: Claussen: Weak calibrator tests (this week)
  • Spectral Line: Momjian: Zeeman followup (this week)
  • Spectral Line: Dhawan: Follow-up on frequency offset phase jumps characterization.
  • Spectral Line: Ott: Doppler tracking sampling follow-up
  • Polarization: Antenna cross polarization (Perley; Need script.); LSCX on 3C147; Postponed until wideband X available on EA24.

RSRO/ECSO Reports

Misc:

Commissioning Group updates

AOB

Misc Actions

  • [Frail/Butler]: Enable restrictions based on proximity to sun; review OST weather trending.
  • [McMullin/Frail]: %REDWrite up disposition of RR items.
  • [McMullin]: Confirm restriction of RSRO scan length at 90 s
  • [Owen]: Outline issue with 90s scan length restriction.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • Development of WIDAR guidelines document (Michael)
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [Rupen]: X, Ku band subbands are inverted (IF 1 is labelled as IF4, IF2 as 3, IF3 as 2, IF 4 as 1). Noted that this had been tested but there were enough changes to warrant a re-do.
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Seen on EA07, EA25 (both RR and LL)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.

-- JosephMcMullin - 2010-11-15
Topic revision: r4 - 2010-11-16, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback